# Crossed products by compact group actions with the weak tracial Rokhlin property

### v2

@@ 3. The weak tracial Rokhlin property with comparison
Proof of sufficiency. Let $\varepsilon>0$, finite set $F\subseteq A$, finite set $S\subseteq C(G)$, $x\in A_{+}$ with $\|x\|=1$, and $y\in(A^{\alpha})_{+}\setminus\{0\}$. Choose the given $d$ and $\varphi$. Choose a contractive positive element $h=(h_{n})_{n\in\mathbb{N}}\in l^{\infty}_{\alpha}(\mathbb{N},A)$ such that $\pi_{A}(h)=d$. By averaging over $G$, we may assume $h$ is $\alpha^{\infty}$-invariant. Then $\pi_{A}$ defines a surjection from $\overline{hl^{\infty}_{\alpha}(\mathbb{N},A)h}$ to $\overline{dA_{\infty,\alpha}d}$. By the Choi-Effros lifting Theorem, we can choose a lift $\theta=(\theta_{n})_{n\in\mathbb{N}}\colon C(G)\to\overline{hl^{\infty}_{\alpha}(\mathbb{N},A)h}$ for $\varphi$ consisting of completely positive contractive maps $\theta_{n}\colon C(G)\to\overline{h_{n}Ah_{n}}$. We claim the following properties hold:
@@
    $\lim_{n\to\infty}\|\theta_{n}(1)\theta_{n}(f_{1}f_{2})-\theta_{n}(f_{1})\theta_{n}(f_{2})\|=0$ for all $f_{1},f_{2}\in C(G)$.
@@
    $\lim_{n\to\infty}\|\theta_{n}(f)a-a\theta_{n}(f)\|=0$ for all $f\in C(G)$ and $a\in F$.
@@
for $g\in G$. Since $\pi_{A}\circ\theta=\varphi$ is equivariant, we have $\lim_{n\to\infty}\rho_{n}(g)=0$ for all $g\in G$, and $\rho_{m}(g)\leq\rho_{n}(g)$ whenever $g\in G$ and $m\geq n$. Let $\sigma\colon l_{\alpha}^{\infty}(\mathbb{N},A)\to l_{\alpha}^{\infty}(\mathbb{N},A)$ be defined by $\sigma((a_{n})_{n\in\mathbb{N}})=(a_{2},a_{3},\dots)$. Now
@@
Since the weak tracial Rokhlin property with comparison holds, by Lemma 3.6, for $S_{n}=\{f_{1},f_{2},\dots,f_{n}\}$, $F_{n}=\{a_{1},a_{2},\dots,a_{n}\}$, $\frac{1}{n}$, $x$, $y$, we obtain contractive positive elements $h_{n}\in A^{\alpha}$ and equivariant completely positive contractive maps $\psi_{n}\colon C(G)\to\overline{h_{n}Ah_{n}}$ such that
@@
Define an $\alpha^{\infty}$-invariant contractive positive element $h\in l^{\infty}_{\alpha}(\mathbb{N},A)$ by $h=(h_{n})_{n\in\mathbb{N}}$. Define an equivariant completely positive contractive map $\psi\colon C(G)\to l^{\infty}(\mathbb{N},A)$ by $\psi(f)=(\psi_{n}(f))_{n\in\mathbb{N}}$ for $f\in C(G)$. Clearly $\psi(1)=h$. By equivariance, the range of $\psi$ is contained in $l^{\infty}_{\alpha}(\mathbb{N},A)$. Set $d=\pi_{A}(h)$ and $\varphi=\pi_{A}\circ\psi$. It is easy to see that $\varphi$ is an equivariant completely positive contractive order zero map from $C(G)$ to $\overline{d(A_{\infty,\alpha}\cap A^{\prime})d}$ with $\varphi(1)=d$.
@@
[17, Lemma 4.2] Let $A$ be a unital C*-algebra, let $G$ be a compact group, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be a continuous action. Let $\varphi\colon C(G,A)\to C(G,A)$ be the averaging process defined above, and let $\gamma\colon G\to\mathrm{Aut}(C(G,A))$ be the diagonal action. Given $\varepsilon>0$ and a compact set $F\subseteq C(G,A)$, there exist $\delta>0$, a finite set $K\subseteq G$, and continuous functions $f_{k}\in C(G)$ for $k\in K$ such that
@@
The following proof references [17, Proposition 4.3].
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action with the weak tracial Rokhlin property with comparison. For every $\varepsilon>0$, every compact set $F\subseteq A$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist $\delta>0$, a finite set $K\subseteq G$, a partition of unity $(f_{k})_{k\in K}\subseteq C(G)$, a contractive positive element $d\in A^{\alpha}$, and a completely positive contractive map $\psi\colon C(G)\to\overline{dAd}$ such that
@@
We now present the most important theorem of this section, which is a key tool for transferring properties from the original algebra to the fixed point algebra. The proof references [19, Theorem 3.1].
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple separable infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action with the weak tracial Rokhlin property with comparison. For every $\varepsilon>0$, every compact set $F_{1}\subseteq A$, every compact set $F_{2}\subseteq A^{\alpha}$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a contractive positive element $d\in A^{\alpha}$ and a completely positive contractive map $\varphi\colon A\to\overline{dA^{\alpha}d}$ such that
@@
Let $\varepsilon_{0}=\frac{\varepsilon}{7}$, take the continuous function $f(t)=t^{\frac{1}{2}}$, $t\in[0,1]$, and apply [5, Lemma 1.5] to $\varepsilon_{0}$ to obtain $\delta_{1}>0$. Let $\delta_{2}=\min(\varepsilon_{0},\ \delta_{1})$. Without loss of generality, we may assume that $\|a\|\leq 1$ for all $a\in F_{1}\cup F_{2}$. For the compact set $F=F_{1}\cup F_{2}$ (without loss of generality, assume it is $\alpha$-invariant), $\delta_{2}$, $x\in A_{+}$ with $\|x\|=1$, and $y\in(A^{\alpha})_{+}\setminus\{0\}$, using Proposition 3.15, we can obtain a positive number $\delta>0$, a finite subset $K\subseteq G$, a partition of unity $(f_{k})_{k\in K}\subseteq C(G)$, a contractive positive element $d\in A^{\alpha}$, and a completely positive contractive map $\psi\colon C(G)\to\overline{dAd}$ such that
@@
For condition (5), for any $n\in\mathbb{N}$, assume $\{S_{n}\}$ is a sequence of finite subsets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F_{2}$, $x$, and $y$, there exist a contractive positive element $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Define a completely positive contractive order zero map $\Phi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\Phi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. We have
@@ 4. Simplicity of the fixed point algebra and the crossed product
For the case of finite groups, it follows from [13, Proposition 3.2] that the weak tracial Rokhlin property implies pointwise outerness. Therefore, by [33, Theorem 3.1], the crossed product is simple. However, without the discreteness of the group, pointwise outerness of the action does not imply simplicity of the crossed product.
@@
Take a nonzero positive element $z\in I$. Let $F=\{x_{i}:i=1,2,\dots,n\}$. Suppose $\{S_{n}\}$ is a sequence of finite sets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $z$, $S_{n}$, and $F$, there exist a contractive positive element $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Define a completely positive contractive order zero map $\varphi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\varphi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. We have
@@
We first prove infinite-dimensionality. We will show that for any $n\in\mathbb{N}$, there exist nonzero mutually orthogonal positive elements $x_{1},\dots,x_{n}\in A^{\alpha}$. Since $A$ is simple and infinite-dimensional, there exist $y_{1},\dots,y_{n}\in A_{+}$ with $\|y_{k}\|=1$ for $k=1,\dots,n$. For given $\frac{1}{3}$, $n$, choose $0<\delta<\frac{1}{3}$ satisfying the condition of [34, Lemma 2.5.12].
By Theorem 3.16, for $\delta<\frac{1}{3}$, $n=2$, $x=y=1$, $F_{1}=\{y_{1},\dots,y_{n}\}$, and $F_{2}=\varnothing$. There exist a contractive positive element $d\in A^{\alpha}$ and a completely positive contractive map $\varphi\colon A\to\overline{dA^{\alpha}d}$ such that
@@
From (1), we have $\|\varphi(y_{j})\varphi(y_{k})\|<\delta$ when $j\neq k$. By [34, Lemma 2.5.12], there exist $x_{1},\dots,x_{n}\in(A^{\alpha})_{+}$ such that $\|x_{k}-\varphi(y_{k})\|<\frac{1}{3}$ for $k=1,\dots,n$ and $x_{j}x_{k}=0$ for $j\neq k$. Moreover, for $k=1,\dots,n$,
@@
[37, Lemma 3.5] In the setting of Definition 4.3, for any finite-dimensional unitary representation $\pi$ of $G$, the set $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ is a closed two-sided ideal of $(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$.
Recall ([40, Definition 7.1.4]) that an action of a compact group $G$ on a C*-algebra $A$ is saturated if a proper imprimitivity bimodule from $A^{\alpha}$ to $A\rtimes_{\alpha}G$ provides a Morita equivalence.
@@
[37, Theorem 3.6] Let $A$ be a C*-algebra, let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action of a compact group on $A$. The following are equivalent:
@@
Let $\pi\colon G\to\mathrm{U}(\mathcal{H}_{\pi})$ belong to $\hat{G}$. By Lemma 4.5, we will show that $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}=(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$. By Lemma 4.5, $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ is a closed two-sided ideal of $(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$. Therefore, it suffices to prove that $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ contains an invertible element.
Let $d_{\pi}=\mathrm{dim}(\mathcal{H}_{\pi})$ and identify $\mathcal{H}_{\pi}$ with $\mathbb{C}^{d_{\pi}}$. Let $\delta=\min(\frac{1}{4d_{\pi}^{3}},\frac{1}{4})$. Endow $C(G)$ with the action $\mathrm{Lt}$. Denote by $\pi^{*}$ the function $g\mapsto\pi(g)^{*}$. A simple calculation shows that $\pi^{*}$ belongs to $C(G)_{2}(\pi)$. Indeed, fix $g\in G$, for $h\in G$, we have
@@
Thus, $(\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\mathrm{Lt}_{g})(\pi^{*})=\pi^{*}(\pi(g)\otimes 1_{C(G)})$, so $\pi^{*}\in C(G)_{2}(\pi)$. For $j,k=1,2,\dots,d_{\pi}$, let $\pi_{j,k}\in C(G)$ be the function $g\mapsto\pi(g)_{j,k}$. Applying Lemma 3.13, we obtain a contractive positive element $r\in(A_{\infty,\alpha}\cap A^{\prime})^{\alpha_{\infty}}$ and an equivariant completely positive contractive order zero map $\varphi\colon C(G)\to\overline{r(A_{\infty,\alpha}\cap A^{\prime})r}$ satisfying $\varphi(1)=r$ and $1-r\precsim_{p.w.-A^{\alpha}}r$. Note that $r$ is $\alpha_{\infty}$-invariant. By Lemma 4.6, it is easy to see that we obtain an equivariant completely positive contractive order zero map $\sqrt{\varphi}$. Then
@@
Since $C(G)$ is amenable and separable, by the Choi–Effros lifting theorem, for $\delta$, $F=\{1_{A}\}$, and $S=\{\pi_{j,k},(\pi_{j,k})^{*}\colon j,k=1,2,\dots,d_{\pi}\}$, we obtain a contractive positive element $g\in A^{\alpha}$ such that $(1-g-\delta)_{+}\precsim_{A^{\alpha}}g$ and an $(S,F,\delta)$-approximately equivariant central strongly order zero completely positive contractive map $\psi\colon C(G)\to\overline{gAg}$ with $\psi(1)=g^{\frac{1}{2}}$ such that for all $f_{1},f_{2}\in S$, $\|g^{\frac{1}{2}}\psi(f_{1}f_{2})-\psi(f_{1})\psi(f_{2})\|<\delta$. As in Lemma 3.6, we will assume that $\psi$ is equivariant. Then
@@
Since $A\rtimes_{\alpha}G$ is simple, $A^{\alpha}$ is a full corner of $A\rtimes_{\alpha}G$ (see [46, Corollary]), hence they are strongly Morita equivalent. Furthermore, since both $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are separable, they are stably isomorphic. ∎

### v3

@@ 3. The weak tracial Rokhlin property with comparison
Proof of sufficiency. Let $\varepsilon>0$, finite set $F\subseteq A$, finite set $S\subseteq C(G)$, $x\in A_{+}$ with $\|x\|=1$, and $y\in(A^{\alpha})_{+}\setminus\{0\}$. Choose the given $d$ and $\varphi$. Choose a positive contraction $h=(h_{n})_{n\in\mathbb{N}}\in l^{\infty}_{\alpha}(\mathbb{N},A)$ such that $\pi_{A}(h)=d$. By averaging over $G$, we may assume $h$ is $\alpha^{\infty}$-invariant. Then $\pi_{A}$ defines a surjection from $\overline{hl^{\infty}_{\alpha}(\mathbb{N},A)h}$ to $\overline{dA_{\infty,\alpha}d}$. By the Choi-Effros lifting Theorem, we can choose a lift $\theta=(\theta_{n})_{n\in\mathbb{N}}\colon C(G)\to\overline{hl^{\infty}_{\alpha}(\mathbb{N},A)h}$ for $\varphi$ consisting of completely positive contractive maps $\theta_{n}\colon C(G)\to\overline{h_{n}Ah_{n}}$. We claim the following properties hold:
@@
    $\lim\limits_{n\to\infty}\|\theta_{n}(1)\theta_{n}(f_{1}f_{2})-\theta_{n}(f_{1})\theta_{n}(f_{2})\|=0$ for all $f_{1},f_{2}\in C(G)$.
@@
    $\lim\limits_{n\to\infty}\|\theta_{n}(f)a-a\theta_{n}(f)\|=0$ for all $f\in C(G)$ and $a\in F$.
@@
for $g\in G$. Since $\pi_{A}\circ\theta=\varphi$ is equivariant, we have $\lim_{n\to\infty}\rho_{n}(g)=0$ for all $g\in G$, and $\rho_{m}(g)\leq\rho_{n}(g)$ whenever $g\in G$ and $m\geq n$. Let $\sigma\colon l_{\alpha}^{\infty}(\mathbb{N},A)\to l_{\alpha}^{\infty}(\mathbb{N},A)$ be the map defined by $\sigma((a_{n})_{n\in\mathbb{N}})=(a_{2},a_{3},\dots)$. Now
@@
Since the weak tracial Rokhlin property with comparison holds, by Lemma 3.6, for $S_{n}=\{f_{1},f_{2},\dots,f_{n}\}$, $F_{n}=\{a_{1},a_{2},\dots,a_{n}\}$, $\frac{1}{n}$, $x$, $y$, we obtain positive contractions $h_{n}\in A^{\alpha}$ and equivariant completely positive contractive maps $\psi_{n}\colon C(G)\to\overline{h_{n}Ah_{n}}$ such that the following hold:
@@
Define an $\alpha^{\infty}$-invariant positive contraction $h\in l^{\infty}_{\alpha}(\mathbb{N},A)$ by $h=(h_{n})_{n\in\mathbb{N}}$. Define an equivariant completely positive contractive map $\psi\colon C(G)\to l^{\infty}(\mathbb{N},A)$ by $\psi(f)=(\psi_{n}(f))_{n\in\mathbb{N}}$ for $f\in C(G)$. Clearly $\psi(1)=h$. By equivariance, the range of $\psi$ is contained in $l^{\infty}_{\alpha}(\mathbb{N},A)$. Set $d=\pi_{A}(h)$ and $\varphi=\pi_{A}\circ\psi$. It is easy to see that $\varphi$ is an equivariant completely positive contractive order zero map from $C(G)$ to $\overline{d(A_{\infty,\alpha}\cap A^{\prime})d}$ with $\varphi(1)=d$.
@@
[21, Lemma 4.2] Let $A$ be a unital C*-algebra, let $G$ be a compact group, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be a continuous action. Let $\varphi\colon C(G,A)\to C(G,A)$ be the averaging process defined above, and let $\gamma\colon G\to\mathrm{Aut}(C(G,A))$ be the diagonal action. Given $\varepsilon>0$ and a compact set $F\subseteq C(G,A)$, there exist $\delta>0$, a finite set $K\subseteq G$, and continuous functions $f_{k}\in C(G)$ for $k\in K$ such that the following hold:
@@
The following proof references [21, Proposition 4.3].
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action with the weak tracial Rokhlin property with comparison. For every $\varepsilon>0$, every compact set $F\subseteq A$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist $\delta>0$, a finite set $K\subseteq G$, a partition of unity $(f_{k})_{k\in K}\subseteq C(G)$, a positive contraction $d\in A^{\alpha}$, and a completely positive contractive map $\psi\colon C(G)\to\overline{dAd}$ such that the following hold:
@@
We now present the most important theorem of this section, which is a key tool for transferring properties from the original algebra to the fixed point algebra. The proof references [22, Theorem 3.1].
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple separable infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action with the weak tracial Rokhlin property with comparison. For every $\varepsilon>0$, every compact set $F_{1}\subseteq A$, every compact set $F_{2}\subseteq A^{\alpha}$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a positive contraction $d\in A^{\alpha}$ and a completely positive contractive map $\varphi\colon A\to\overline{dA^{\alpha}d}$ such that the following hold:
@@
Let $\varepsilon_{0}=\frac{\varepsilon}{7}$, take the continuous function $f(t)=t^{\frac{1}{2}}$, $t\in[0,1]$, and apply [5, Lemma 1.5] to $\varepsilon_{0}$ to obtain $\delta_{1}>0$. Let $\delta_{2}=\min(\varepsilon_{0},\ \delta_{1})$. Without loss of generality, we may assume that $\|a\|\leq 1$ for all $a\in F_{1}\cup F_{2}$. For the compact set $F=F_{1}\cup F_{2}$ (without loss of generality, assume it is $\alpha$-invariant), $\delta_{2}$, $x\in A_{+}$ with $\|x\|=1$, and $y\in(A^{\alpha})_{+}\setminus\{0\}$, using Proposition 3.15, we can obtain a positive number $\delta>0$, a finite subset $K\subseteq G$, a partition of unity $(f_{k})_{k\in K}\subseteq C(G)$, a positive contraction $d\in A^{\alpha}$, and a completely positive contractive map $\psi\colon C(G)\to\overline{dAd}$ such that the following hold:
@@
For condition (5), for any $n\in\mathbb{N}$, assume $\{S_{n}\}$ is a sequence of finite subsets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $S_{n}$, $F_{2}$, $x$, and $y$, there exist a positive contraction $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that the following hold:
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. Define a completely positive contractive order zero map $\Phi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\Phi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. We have the following hold:
@@ 4. Simplicity of the fixed point algebra and the crossed product
For the case of finite groups, it follows from [13, Proposition 3.2] that the weak tracial Rokhlin property implies pointwise outerness. Therefore, by [32, Theorem 3.1], the crossed product is simple. However, without the discreteness of the group, pointwise outerness of the action does not imply simplicity of the crossed product.
@@
Take a nonzero positive element $z\in I$. Let $F=\{x_{i}:i=1,2,\dots,n\}$. Suppose $\{S_{n}\}$ is a sequence of finite sets in $A$ such that $\cup_{n\in\mathbb{N}}S_{n}$ is dense in $A$ and $F\subseteq S_{n}$ for all $n\in\mathbb{N}$. For $\frac{1}{n}$, $z$, $S_{n}$, and $F$, there exist a positive contraction $d_{n}\in A^{\alpha}$ and a completely positive contractive map $\varphi_{n}\colon A\to\overline{d_{n}A^{\alpha}d_{n}}$ such that the following hold:
@@
Let $\pi_{A^{\alpha}}\colon l^{\infty}(\mathbb{N},A^{\alpha})\to(A^{\alpha})_{\infty}$ be the quotient map. Denote $d=(d_{n})_{n=1}^{\infty}\in l^{\infty}(\mathbb{N},A^{\alpha})$ and $r=\pi_{A^{\alpha}}(d)$. Define a completely positive contractive order zero map $\varphi\colon A\to\overline{r(A^{\alpha})_{\infty}r}$ by $\varphi(a)=\pi_{A^{\alpha}}(\{\varphi_{1}(a),\varphi_{2}(a),\dots\})$ for all $a\in A$. We have the following hold:
@@
We first prove infinite-dimensionality. We will show that for any $n\in\mathbb{N}$, there exist nonzero mutually orthogonal positive elements $x_{1},\dots,x_{n}\in A^{\alpha}$. Since $A$ is simple and infinite-dimensional, there exist $y_{1},\dots,y_{n}\in A_{+}$ with $\|y_{k}\|=1$ for $k=1,\dots,n$. For given $\frac{1}{3}$, $n$, choose $0<\delta<\frac{1}{3}$ satisfying the condition of [33, Lemma 2.5.12].
By Theorem 3.16, for $\delta<\frac{1}{3}$, $n=2$, $x=y=1$, $F_{1}=\{y_{1},\dots,y_{n}\}$, and $F_{2}=\varnothing$. There exist a positive contraction $d\in A^{\alpha}$ and a completely positive contractive map $\varphi\colon A\to\overline{dA^{\alpha}d}$ such that the following hold:
@@
From (1), we have $\|\varphi(y_{j})\varphi(y_{k})\|<\delta$ when $j\neq k$. By [33, Lemma 2.5.12], there exist $x_{1},\dots,x_{n}\in(A^{\alpha})_{+}$ such that $\|x_{k}-\varphi(y_{k})\|<\frac{1}{3}$ for $k=1,\dots,n$ and $x_{j}x_{k}=0$ for $j\neq k$. Moreover, for $k=1,\dots,n$,
@@
[36, Lemma 3.5] In the setting of Definition 4.3, for any finite-dimensional unitary representation $\pi$ of $G$, the set $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ is a closed two-sided ideal of $(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$.
Recall ([39, Definition 7.1.4]) that an action of a compact group $G$ on a C*-algebra $A$ is saturated if a suitable completion of $A$ is a Morita equivalence bimodule from $A^{\alpha}$ to $A\rtimes_{\alpha}G$. The following theorem gives us some essentially known results on saturation, the strong Arveson spectrum, and simplicity.
@@
[36, Theorem 3.6] Let $A$ be a C*-algebra, let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action of a compact group on $A$. The following are equivalent:
@@
Let $\pi\colon G\to\mathrm{U}(\mathcal{H}_{\pi})$ belong to $\hat{G}$. By Theorem 4.5, we will show that $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}=(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$. By Lemma 4.4, $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ is a closed two-sided ideal of $(B(\mathcal{H}_{\pi})\otimes A)^{\mathrm{Ad}(\pi)\otimes\alpha}$. Therefore, it suffices to prove that $\overline{A_{2}(\pi)^{*}A_{2}(\pi)}$ contains an invertible element.
It follows from the Peter–Weyl theorem that all irreducible unitary representations of a compact group are finite-dimensional. Let $d_{\pi}=\mathrm{dim}(\mathcal{H}_{\pi})$ and identify $\mathcal{H}_{\pi}$ with $\mathbb{C}^{d_{\pi}}$. Let $\delta=\min(\frac{1}{4d_{\pi}^{3}},\frac{1}{4})$. Endow $C(G)$ with the action $\mathrm{Lt}$. Denote by $\pi^{*}$ the function $g\mapsto\pi(g)^{*}$. A simple calculation shows that $\pi^{*}$ belongs to $C(G)_{2}(\pi)$. Indeed, fix $g\in G$, for $h\in G$, we have
@@
Thus, $(\mathrm{id}_{B(\mathcal{H}_{\pi})}\otimes\mathrm{Lt}_{g})(\pi^{*})=\pi^{*}(\pi(g)\otimes 1_{C(G)})$, so $\pi^{*}\in C(G)_{2}(\pi)$. For $j,k=1,2,\dots,d_{\pi}$, let $\pi_{j,k}\in C(G)$ be the function $g\mapsto\pi(g)_{j,k}$. Applying Lemma 3.13, we obtain a positive contraction $r\in(A_{\infty,\alpha}\cap A^{\prime})^{\alpha_{\infty}}$ and an equivariant completely positive contractive order zero map $\varphi\colon C(G)\to\overline{r(A_{\infty,\alpha}\cap A^{\prime})r}$ satisfying $\varphi(1)=r$ and $1-r\precsim_{p.w.-A^{\alpha}}r$. Note that $r$ is $\alpha_{\infty}$-invariant. By Lemma 4.6, it is easy to see that we obtain an equivariant completely positive contractive order zero map $\sqrt{\varphi}$. Then
@@
Since $C(G)$ is amenable and separable, by the Choi–Effros lifting theorem, for $\delta$, $F=\{1_{A}\}$, and $S=\{\pi_{j,k},(\pi_{j,k})^{*}\colon j,k=1,2,\dots,d_{\pi}\}$, and $x=y=1$, we obtain a positive contraction $g\in A^{\alpha}$ such that $(1-g-\delta)_{+}\precsim_{A^{\alpha}}g$ and an $(S,F,\delta)$-approximately equivariant central strongly order zero completely positive contractive map $\psi\colon C(G)\to\overline{gAg}$ with $\psi(1)=g^{\frac{1}{2}}$ such that for all $f_{1},f_{2}\in S$, $\|g^{\frac{1}{2}}\psi(f_{1}f_{2})-\psi(f_{1})\psi(f_{2})\|<\delta$. As in Lemma 3.6, we will assume that $\psi$ is equivariant. Then
@@
Since $A\rtimes_{\alpha}G$ is simple, $A^{\alpha}$ is a full corner of $A\rtimes_{\alpha}G$ (see [45, Corollary]), hence they are strongly Morita equivalent. Furthermore, since both $A^{\alpha}$ and $A\rtimes_{\alpha}G$ are separable, they are stably isomorphic. ∎